A wooden bench 100cm long and a mass of 95g can be balanced from a knife edge when 5g mass is hung 10cm from one end,how far is the knife edge from the center of the edge


1
Expert's answer
2021-07-07T12:53:01-0400

The bench is in equilibrium if


"\tau_1 = \tau_2"

where "\tau_1, \tau_2" are the magnitude of the torques that rotate the bench clockwise and counterclockwise about the knife edge respectively.

By definition:


"\tau_1 = m_1gl_1\\
\tau_2 = m_2gl_2"

where "m_1 = 5g, m_2 = 95g, l_1 = 50cm-5cm-l_2 = 45cm-l_2" is the distance of the 5g mass from the knife edge, "l_2" is the distance of the knife edge from the center of the bench, and "g" is the gravitational acceleration.

Thus, obtain:

"5\cdot g\cdot (45-l_2) = 95\cdot g\cdot l_2\\
5\cdot (45-l_2) = 95\cdot g\cdot l_2\\
225 - 5l_2 = 95l_2\\
l_2 = 2.5cm"

Answer. 2.5 cm.